# Graph-based Approaches and Functionalities in Retrieval-Augmented Generation: A Comprehensive Survey

### v1

@@ Abstract.
Large language models (LLMs) struggle with the factual error during inference due to the lack of sufficient training data and the most updated knowledge, leading to the hallucination problem. Retrieval-Augmented Generation (RAG) has gained attention as a promising solution to address the limitation of LLMs, by retrieving relevant information from external source to generate more accurate answers to the questions. Given the pervasive presence of structured knowledge in the external source, considerable strides in RAG have been made to employ the techniques related to graphs and achieve more complex reasoning based on the topological information between knowledge entities. However, there is currently neither unified review examining the diverse roles of graphs in RAG, nor a comprehensive resource to help researchers navigate and contribute to this evolving field. This survey offers a novel perspective on the functionality of graphs within RAG and their impact on enhancing performance across a wide range of graph-structured data. It provides a detailed breakdown of the roles that graphs play in RAG, covering database construction, algorithms, pipelines, and tasks. Finally, it identifies current challenges and outline future research directions, aiming to inspire further developments in this field. Our graph-centered analysis highlights the commonalities and differences in existing methods, setting the stage for future researchers in areas such as graph learning, database systems, and natural language processing.
@@
ccs: Information systems Information retrievalccs: Computing methodologies Natural language processing
@@ 1. Introduction
Large Language Models (LLMs) have demonstrated high effectiveness in a wide spectrum of human-centric tasks, including text summarization (ding2024evaluation, ), question answering (ref:ewek-qa, ), machine translation (feng2024improving, ), and code generation (liu2024your, ). Early models like BERT (ref:bert, ) and RoBERTa (ref:roberta, ) introduced advanced techniques for leveraging contextual embeddings, significantly enhancing performance of different natural language understanding tasks. Inspired by these fundamental works, models such as GPT-4 (ref:gpt4, ), LLaMA 3 (ref:llama, ), and more recently Deepseek-R1 (ref:deepseek, ), have set new benchmarks by utilizing transformer decoder-only architectures and scaling up model sizes. These LLMs excel at generating human-like responses, adapting to a wide range of domains, and supporting complex reasoning. All of these advancements have led to a growing consensus that LLMs represent a key pathway to artificial general intelligence (AGI) (bubeck2023sparks, ).
@@
Despite these successes, LLMs face two major challenges: they are prone to factual errors and hallucinations, and they frequently struggle with real-world knowledge that inherently represents structured information (lavrinovics2025knowledge, ; agrawal2023can, ; zhu2024kg, ). The first challenge arises because the retraining of LLMs to incorporate new or domain-specific knowledge is computationally expensive (ref:talk, ), leaving models susceptible to producing logically coherent but factually incorrect inferences, which is also known as the hallucination problem (Shuster2021retrieval, ). For the second challenge, knowledge graphs like DBpedia (dbpedia, ), Wikidata (ref:wikidata, ), and Freebase (ref:freebase, ) are primary examples of how entities and relationships are organized in a structured format, yet existing LLM-based approaches often fail to fully exploit these relational patterns, especially for complex or multi-hop reasoning tasks.
@@
These two challenges call for retrieval mechanisms that enable LLMs to dynamically access external and structured knowledge during inference. Retrieval-Augmented Generation (RAG) helps solve the first challenges by retrieving relevant external information, thereby reducing hallucinations and enhancing factual accuracy (ref:ragsurvey5, ; ref:ragsurvey3, ). With respect to the second challenge, the integration of graph data management with RAG further enhances these capabilities by effectively organizing knowledge into organized forms such as knowledge graphs (ref:GPT4Graph, ). This integration captures complex relational patterns and supports multi-hop reasoning tasks such as node classification (ref:Graph-LLM, ; ref:graphtranslator, ), link prediction (ref:wn, ; ref:kg-r3, ), and subgraph extraction (ref:hipporag, ; ref:rasr, ; ref:dalk, ). Consequently, graph-enhanced RAG provides a powerful approach to improving LLM’s factual reliability and their ability to execute complex, structured reasoning tasks.
@@
The application of graph techniques in RAG has rapidly expanded, leading several recent surveys on LLMs to discuss graph-based methods, such as Peng’s 2024 survey (peng2024graph, ) and Zhang et al.’s more recent work in 2025 (zhang2025survey, ). Nevertheless, existing surveys primarily focus on general RAG architectures or applications, without placing emphasis on the fundamental graph techniques underlying RAG or offering a comprehensive review from the perspective of graph data management. In contrast, this paper fills this notable gap by providing a meticulous and detailed decomposition of the RAG framework, thoroughly examining the concrete graph techniques applied in representative works. We explicitly discuss the rationale behind integrating these graph methods into RAG systems and highlight their specific advantages in enhancing answer accuracy and efficiency. Furthermore, based on insights drawn from over 200 recent RAG studies, we systematically categorize the use of graph techniques across the core components of RAG—including database construction, retrieval or prompting algorithms, processing pipelines, and task scenarios. Unlike previous surveys (peng2024graph, ; ref:ragsurvey1, ; ref:ragsurvey2, ; ref:ragsurvey5, ), this work offers an in-depth exploration from the viewpoint of graph data management with respect to database construction, query algorithms, prompt structure, and pipeline design, presenting novel insights into how detailed graph methods can significantly drive advancements in RAG research. The contributions of this work are summarized as follows:
@@ 2. Comparison with Existing Surveys
Surveys of LLMs. A large number of surveys (yang2023harnessing, ; huang2022towards, ; naveed2023comprehensive, ; zhao2023survey, ; minaee2024large, ; hadi2023survey, ; ref:llm_survey, ; ref:cot-survey, ) have explored LLMs from various perspectives over the years, reflecting their growing prominence in AI research. Early works (yang2023harnessing, ; hadi2023survey, ) provide comprehensive overviews of foundational models like BERT (ref:bert, ) and GPT-3 (ref:gpt3, ), detailing their architectures, design principles, and advancements compared to earlier models. These surveys emphasize the evolution of LLMs and their capacity to handle zero-shot and few-shot learning tasks, marking a shift from task-specific fine-tuning. Other studies (zhao2023survey, ; minaee2024large, ; shen2023large, ) focus on specific technical aspects, such as pretraining methods, fine-tuning strategies, and alignment techniques for generating reliable outputs. They also address challenges like mitigating biases, ensuring factual accuracy, and aligning outputs with user intent. In addition, domain-specific surveys highlight the applications of LLMs across various fields. Finance-related studies (ref:finalcialgpt2, ; ref:financialgpt, ) explore their use in forecasting and risk analysis, while education-focused works (ref:edu2, ; ref:edu1, ) examine their role in personalized learning and automated assessment. Similarly, healthcare surveys (ref:med2, ; ref:med1, ; ref:med3, ) discuss their contributions to medical diagnosis, patient care, and drug discovery. These surveys collectively showcase the versatility of LLMs across diverse domains and the challenges in their real-world adoption. However, these surveys primarily focus on the general development of LLMs, making them less relevant to the specific scope of our work.
@@
Surveys of RAG. Building on the foundation of LLM surveys, an increasing number of works have concentrated on RAG, exploring its workflow and the integration of external knowledge to enhance generative accuracy. Some surveys (ref:ragsurvey1, ; ref:ragsurvey2, ; ref:ragsurvey3, ; ref:ragsurvey5, ; ref:ragsurvey6, ; ref:ragsurvey7, ; llmir, ) delve into the architecture of RAG, detailing how retrieval mechanisms interact with generative models to create a seamless pipeline for producing context-aware responses. These studies often highlight the modularity of RAG workflows, emphasizing the flexibility of incorporating different retrieval strategies, such as vector-based retrieval and hybrid methods, to optimize performance across diverse tasks. Other surveys (tonmoy2024comprehensive, ; ref:hallucination, ; rawte2023survey, ; zhang2023siren, ) prioritize addressing the hallucination problem, analyzing how external information retrieval mitigates factual inaccuracies, and ensures the generated responses align closely with real-world knowledge. They discuss the effectiveness of external retrieval in outputs of grounding models and reducing the risk of plausible but incorrect information generation. Additionally, some works (zhao2024retrieval, ; ref:ragsurvey4, ; zhou2024trustworthiness, ) explore the broader applications of RAG, demonstrating its utility in tasks like question answering, summarization, and domain-specific knowledge generation. However, given the importance of graphs as a fundamental structure in RAG, graph-related applications are not a central focus of these works.
@@
Surveys of LLMs on graphs. Existing surveys of LLMs on graphs can broadly be categorized into two categories. The first category focuses on how LLMs can enhance prediction tasks on graphs by utilizing their ability to comprehend natural language (ref:gfmsurvey1, ; ref:gfmsurvey2, ; ref:gfmsurvey3, ; ref:gfmsurvey4, ; ref:gfmsurvey7, ). These surveys examine techniques such as (i) feature augmentation, where LLMs enhance node or edge attributes; (ii) feature alignment, which bridges textual and graph representations; and (iii) structural enhancement, leveraging LLMs to model intricate graph relationships and topological patterns. These works emphasize the potential of LLMs in boosting the performance of tasks like node classification, link prediction, and graph-based recommendation systems. Surveys in another category (ref:gfmsurvey5, ; ref:gfmsurvey6, ; ref:gfmsurvey8, ) investigate how LLMs enhance traditional symbolic knowledge bases through three key approaches: (i) deploying LLMs as knowledge graph builders and controllers; (ii) leveraging structured knowledge to improve LLM pretraining; and (iii) enabling LLM-augmented symbolic reasoning to refine logical inference. With the approaches above, these methods showcase the potential of LLMs to enrich and expand the capabilities of knowledge graphs in both construction and reasoning tasks. Another important related work (peng2024graph, ) examines the general workflow of how RAG leverages external knowledge graphs to enhance LLM predictions. While their work provides a valuable overview of the general RAG workflow, it has two critical limitations: (i) it focuses on surface-level interactions between LLMs and graph-related tasks, overlooking the diverse roles that graphs play in advancing RAG systems (e.g., structural reasoning, and knowledge alignment); and (ii) it lacks actionable guidance for graph researchers seeking to apply their expertise to LLM-driven methodologies.
In contrast, our work bridges this gap by systematically analyzing the multifaceted roles of graphs in RAG-driven systems, from the angle of graph data management including database construction, query algorithms, prompt structure, pipeline design and downstream tasks. As a resource tailored for graph researchers exploring LLM applications, we not only delineate how graphs enhance LLM capabilities but also provide actionable strategies for leveraging graph expertise to advance RAG systems, fostering interdisciplinary synergy.
@@ 3.2. RAG with Graph Data Management
Graphs, with their inherent ability to model complex relationships and dependencies, serve as powerful tools for structuring external knowledge and enabling sophisticated reasoning in RAG workflows. Existing literature (ref:gfmsurvey7, ; peng2024graph, ) primarily focuses on how LLMs enhance graph performance or provide a brief introduction of the RAG process involving graphs. However, several critical questions remain unanswered: What benefits does the graph structure offer to RAG? Which graph patterns can enhance the effectiveness or efficiency of LLMs responses? What future directions should be pursued to develop graph algorithms that synergize with LLMs? To address these compelling questions, we approach this survey explicitly from the perspective of graph data management, systematically investigating how graph techniques enhance each core stage of the RAG process, as illustrated in Figure 2. Specifically, we highlight how graph data management methods significantly (i) improve knowledge base construction; (ii) optimize retrieval and prompting algorithms through graph-based indexing, querying, and reasoning; (iii) streamline data processing pipelines via graph-structured workflows; and (iv) effectively support graph-oriented tasks.
@@ 4.1.1. A Unified View
In this section, we focus on existing knowledge graphs, which play a crucial role in many RAG systems by serving as structured repositories of factual knowledge. To enable efficient querying and traversal of entities and their relationships, many existing methods (ref:grag, ; ref:kaping, ; ref:unikgqa, ) directly retrieve knowledge relevant to the query from well-established knowledge graphs such as FreeBase (ref:freebase, ), T-REx (ref:t-rex, ), and WebQSP (ref:webqsp, ). These widely used databases offer comprehensive scopes for knowledge retrieval, encompassing a vast array of entities and relationships across diverse domains. For instance, T-REx (ref:t-rex, ) provides extensive alignments between natural language expressions and knowledge base triples, facilitating the integration of structured data with textual information. By organizing nodes and edges along with their textual attributes, these structured repositories allow RAG systems to efficiently access additional topological information, supporting advanced operations such as entity retrieval, relationship discovery, and complex multi-hop queries.
@@ 4.1.2. Existing Knowledge Graphs
Existing KGs Broad coverage, reliable quality General knowledge, less adaptable BioRED (ref:biored, ), QALD-9-plus (ref:qald, ), OpenbookQA (ref:openbookqa, ), CREAK (ref:creak, ), TriviaQA (ref:triviaqa, ), HotpotQA (ref:hotpotqa, ), Mintaka (ref:mintaka, ), MedQA (ref:medqa, ), TUDataset (ref:tudataset, ), CWQ (ref:cwq, ), Beyond I.I.D.  (ref:grailqa, ), CommonsenseQA (ref:commonsenseqa, ), SocialIQA (ref:socialIQA, ), PIQA (ref:piqa, ), RiddleSense (ref:riddle, ), Freebase (ref:freebaseqa, ), ATOMIC (ref:atomic, ), FactKG (ref:factkg, ), MultiHop-RAG (tang2024multihop, ), T-REx (ref:t-rex, ), DBpedia (ref:dbpedia, ), Yago (ref:yago, )
@@
KGs generated from texts Domain-specific, easily updated LM-dependent, computationally intensive GraphRAG (ref:graphrag, ), GRBK (ref:noname5, ), ATLANTIC (ref:atlantic, ), GNN-Ret (ref:gnn-ret, ), HippoRAG (ref:hipporag, ), DALK (ref:dalk, ), KGP (ref:kgp, ), OpenSCR (ref:opencsr, ), MindMap (ref:mindmap, ), FABULA (ref:fabula, ), GER (ref:ger, ), FoodGPT (ref:foodgpt, ), ChatKBQA (ref:chatkbqa, ), MultiQG (ref:noname8, ), HSGE (ref:hsge, ), ReTraCk (ref:retrack, ), RNG-KBQA (ref:rng, ), ArcaneQA (ref:arcaneqa, ), HybridRAG (ref:hybridrag, ), EWEK-QA (ref:ewek-qa, ), KG-FiD (ref:kgfid, ), REANO (ref:reano, ), MedGraphRAG (ref:medgraphrag, ), MINERVA (ref:minerva, )
@@
Freebase (ref:freebase, ) is a large-scale, structured knowledge database designed to organize and store general human knowledge. It combines the efficiency and scalability of structured databases with the rich semantic representation found in sources like Wikipedia (bridge2001wikipedia, ), enabling flexible integration of diverse knowledge. By structuring information into tuples, Freebase facilitates efficient querying, entity linking, and knowledge retrieval across multiple domains. Freebase features an HTTP-based API powered by the Metaweb Query Language (MQL) (flanagan2007developing, ), which facilitates intuitive object-oriented queries and schema evolution. Its complete normalization philosophy ensures unique global identifiers (GUIDs) for real-world entities, while its lightweight typing system supports multiple data. In the context of RAG, Freebase serves as a powerful auxiliary database, enabling methods to retrieve structured knowledge for enhanced reasoning and answer generation. With its ability to handle vast datasets and diverse relationships, Freebase provides a robust foundation for integrating external knowledge into RAG systems, improving their capability to process complex queries and generate context-aware responses.
@@
T-REx (ref:t-rex, ) is a large-scale dataset designed to align natural language from Wikipedia abstracts with knowledge base (KB) triples from Wikidata. Addressing the alignment problem, T-REx connects unstructured text with structured knowledge representations by providing 11 million triples aligned with over 3 million Wikipedia abstracts, covering more than 600 unique predicates. Its customizable alignment pipeline incorporates techniques like predicate linking, co-reference resolution, and distant supervision to create high-quality alignments. Compared to existing datasets, T-REx provides a larger scale, broader predicate coverage, and high accuracy. This dataset is instrumental in advancing tasks like relation extraction, KB population, and question answering, and serves as a foundational resource for RAG systems to retrieve structured knowledge and improve alignment for complex reasoning.

### v2

@@ Abstract.
Large language models (LLMs) struggle with the factual error during inference due to the lack of sufficient training data and the most updated knowledge, leading to the hallucination problem. Retrieval-Augmented Generation (RAG) has gained attention as a promising solution to address the limitation of LLMs, by retrieving relevant information from external source to generate more accurate answers to the questions. Given the pervasive presence of structured knowledge in the external source, considerable strides in RAG have been made to employ the techniques related to graphs and achieve more complex reasoning based on the topological information between knowledge entities. Although recent surveys summarize RAG pipelines, they largely restrict the role of graphs to knowledge-graph traversal, leaving the broader impacts of graph structure on retrieval, prompting, and pipeline control underexplored. This survey offers a novel perspective on the functionality of graphs within RAG and their impact on enhancing performance across a wide range of graph-structured data. It provides a detailed breakdown of the roles that graphs play in RAG, covering database construction, algorithms, pipelines, and tasks. Finally, it identifies current challenges and outline future research directions, aiming to inspire further developments in this field. Our graph-centered analysis highlights the commonalities and differences in existing methods, setting the stage for future researchers in areas such as graph learning, database systems, and natural language processing.
@@
$c$s: Information systems Information retrievalccs: Computing methodologies Natural language processing
@@ 1. Introduction
Large Language Models (LLMs) have demonstrated high effectiveness in a wide spectrum of human-centric tasks, including text summarization (Ding et al., 2024), question answering (Dehghan et al., 2024), machine translation (Feng et al., 2024), and code generation (Liu et al., 2024c). Early models like BERT (Devlin et al., 2019) and RoBERTa (Liu et al., 2019) introduced advanced techniques for leveraging contextual embeddings, significantly enhancing performance of different natural language understanding tasks. Inspired by these fundamental works, models such as GPT-4 (OpenAI, 2024), LLaMA 2 (Touvron et al., 2023), and more recently Deepseek-R1 (DeepSeek-AI et al., 2025), have set new benchmarks by utilizing transformer decoder-only architectures and scaling up model sizes. These LLMs excel at generating human-like responses, adapting to a wide range of domains, and supporting complex reasoning. All of these advancements have led to a growing consensus that LLMs represent a key pathway to artificial general intelligence (AGI) (Bubeck et al., 2023).
@@
Despite these successes, LLMs face two major challenges: they are prone to factual errors and hallucinations, and they frequently struggle with real-world knowledge that inherently represents structured information (Lavrinovics et al., 2025; Agrawal et al., 2023; Zhu et al., 2024a). The first challenge arises because the retraining of LLMs to incorporate new or domain-specific knowledge is computationally expensive (Fatemi et al., 2023), leaving models susceptible to producing logically coherent but factually incorrect inferences, which is also known as the hallucination problem (Shuster et al., 2021). For the second challenge, knowledge graphs like DBpedia (Auer et al., 2007b), Wikidata (Vrandečić and Krötzsch, 2014), and Freebase (Bollacker et al., 2008) are primary examples of how entities and relationships are organized in a structured format, yet existing LLM-based approaches often fail to fully exploit these relational patterns, especially for complex or multi-hop reasoning tasks.
@@
These two challenges call for retrieval mechanisms that enable LLMs to dynamically access external and structured knowledge during inference. Retrieval-Augmented Generation (RAG) helps solve the first challenges by retrieving relevant external information, thereby reducing hallucinations and enhancing factual accuracy (Gao et al., 2024c; Huang and Huang, 2024). With respect to the second challenge, the integration of graph data management with RAG further enhances these capabilities by effectively organizing knowledge into organized forms such as knowledge graphs (Guo et al., 2023). This integration captures complex relational patterns and supports multi-hop reasoning tasks such as node classification (Chen et al., 2024c; Zhang et al., 2024b), link prediction (Dettmers et al., 2018; Pahuja et al., 2023), and subgraph extraction (Gutiérrez et al., 2024; Luo et al., 2023; Li et al., 2024a). Consequently, graph-enhanced RAG provides a powerful approach to improving LLM’s factual reliability and their ability to execute complex, structured reasoning tasks.
@@
The application of graph techniques in RAG has rapidly expanded, leading several recent surveys on LLMs to discuss graph-based methods, such as Han’s survey, Peng’s survey in 2024 (Peng et al., 2024) and Zhang et al.’s more recent work in 2025 (Zhang et al., 2025). While prior surveys provide a useful synthesis of RAG architectures and applications with graphs, they are largely pipeline-centric and often conflate graph-based RAG with knowledge-graph traversal. Consequently, they devote limited attention to the underlying graph techniques, and provide mainly high-level rather than task-specific guidance for applying graph techniques in LLM pipelines. In contrast, this paper fills this notable gap by providing a meticulous and detailed decomposition of the RAG framework, broadening the definition of graph-based RAG and thoroughly examining the concrete graph techniques applied in representative works. We explicitly discuss the rationale behind integrating these graph methods into RAG systems and highlight their specific advantages in enhancing answer accuracy and efficiency. Furthermore, based on insights drawn from over 200 recent RAG studies, we systematically categorize the use of graph techniques across the core components of RAG—including database construction, retrieval or prompting algorithms, processing pipelines, and task scenarios. Unlike previous surveys (Peng et al., 2024; Fan et al., 2024a; Hu and Lu, 2024; Gao et al., 2024c), this work offers an in-depth exploration from the viewpoint of graph data management with respect to database construction, query algorithms, prompt structure, and pipeline design, presenting novel insights into how detailed graph methods can significantly drive advancements in RAG research. The contributions of this work are summarized as follows:
@@ 2. Comparison with Existing Surveys
Surveys of LLMs. A large number of surveys (Yang et al., 2023; Huang and Chang, 2022; Naveed et al., 2023; Zhao et al., 2023b; Minaee et al., 2024; Hadi et al., 2023; Myers et al., 2024; Besta et al., 2024b) have explored LLMs from various perspectives over the years, reflecting their growing prominence in AI research. Early works (Yang et al., 2023; Hadi et al., 2023) provide comprehensive overviews of foundational models like BERT (Devlin et al., 2019) and GPT-3 (Brown et al., 2020), detailing their architectures, design principles, and advancements compared to earlier models. These surveys emphasize the evolution of LLMs and their capacity to handle zero-shot and few-shot learning tasks, marking a shift from task-specific fine-tuning. Other studies (Zhao et al., 2023b; Minaee et al., 2024; Shen et al., 2023) focus on specific technical aspects, such as pretraining methods, fine-tuning strategies, and alignment techniques for generating reliable outputs. They also address challenges like mitigating biases, ensuring factual accuracy, and aligning outputs with user intent.
@@
Surveys of RAG. Building on the foundation of LLM surveys, an increasing number of works have concentrated on RAG, exploring its workflow and the integration of external knowledge to enhance generative accuracy. Some surveys (Fan et al., 2024a; Hu and Lu, 2024; Huang and Huang, 2024; Gao et al., 2024c; Yu et al., 2024; Wu et al., 2024d; Zhu et al., 2023) delve into the architecture of RAG, detailing how retrieval mechanisms interact with generative models to create a seamless pipeline for producing context-aware responses. These studies often highlight the modularity of RAG workflows, emphasizing the flexibility of incorporating different retrieval strategies, such as vector-based retrieval and hybrid methods, to optimize performance across diverse tasks. Other surveys (Tonmoy et al., 2024; Huang et al., 2023a; Rawte et al., 2023; Zhang et al., 2023) prioritize addressing the hallucination problem, analyzing how external information retrieval mitigates factual inaccuracies, and ensures the generated responses align closely with real-world knowledge. They discuss the effectiveness of external retrieval in outputs of grounding models and reducing the risk of plausible but incorrect information generation. Additionally, some works (Zhao et al., 2024b, a; Zhou et al., 2024b) explore the broader applications of RAG, demonstrating its utility in tasks like question answering, summarization, and domain-specific knowledge generation. However, given the importance of graphs as a fundamental structure in RAG, graph-related applications are not a central focus of these works.
@@
Surveys of LLMs on graphs. Existing surveys of LLMs on graphs can broadly be categorized into two categories. The first category focuses on how LLMs can enhance prediction tasks on graphs by utilizing their ability to comprehend natural language (Fan et al., 2024b; Mao et al., 2024; Liu et al., 2024d; Li et al., 2024c; Jin et al., 2024a). These surveys examine techniques such as (i) feature augmentation, where LLMs enhance node or edge attributes; (ii) feature alignment, which bridges textual and graph representations; and (iii) structural enhancement, leveraging LLMs to model intricate graph relationships and topological patterns. These works emphasize the potential of LLMs in boosting the performance of tasks like node classification, link prediction, and graph-based recommendation systems. Surveys in another category (Pan et al., 2023, 2024; Chen, 2024) investigate how LLMs enhance traditional symbolic knowledge bases through three key approaches: (i) deploying LLMs as knowledge graph builders and controllers; (ii) leveraging structured knowledge to improve LLM pretraining; and (iii) enabling LLM-augmented symbolic reasoning to refine logical inference. With the approaches above, these methods showcase the potential of LLMs to enrich and expand the capabilities of knowledge graphs in both construction and reasoning tasks. Another important related work (Peng et al., 2024) examines the general workflow of how RAG leverages external knowledge graphs to enhance LLM predictions.
Comparison with existing surveys on graph-based RAG. Existing surveys on graph-based RAG (Han et al., 2024; Peng et al., 2024) provide a valuable synthesis of the RAG workflow and have helped stabilize terminology and component boundaries. At the same time, they are largely component-centric and tend to equate graph-based RAG with knowledge-graph traversal, which narrows the view of what graphs contribute to retrieval and reasoning. This perspective gives limited attention to analysis grounded in data-management practice and offers only modest, task-level guidance for graph researchers who wish to apply graph techniques in LLM-centric pipelines.
Our survey addresses these gaps along three axes while building on the strengths of prior work. First, it adopts a broader definition of graph-based RAG, encompassing graph-powered databases beyond curated knowledge graphs and clarifying how structural signals inform retrieval and prompting. Second, it provides a deeper analysis from the perspective of graph data management, treating retrieval, prompting, and pipeline topology as design problems with explicit trade-offs in accuracy, latency, token budget, freshness, and provenance. Third, it serves as a comprehensible resource for graph researchers, offering task-wise guidance and decision criteria that indicate when graph structure materially improves outcomes and how to apply it in practice. Together, these contributions complement existing surveys by shifting the emphasis from pipeline components to the roles graphs play in RAG and by translating those roles into actionable methodology.
@@ 3.2. RAG with Graph Data Management
Graphs, with their inherent ability to model complex relationships and dependencies, serve as powerful tools for structuring external knowledge and enabling sophisticated reasoning in RAG workflows. Existing literature (Jin et al., 2024a; Peng et al., 2024) primarily focuses on how LLMs enhance graph performance or provide a brief introduction of the RAG process involving graphs. However, several critical questions remain unanswered: What benefits does the graph structure offer to RAG? Which graph patterns can enhance the effectiveness or efficiency of LLMs responses? What future directions should be pursued to develop graph algorithms that synergize with LLMs? To address these compelling questions, we approach this survey explicitly from the perspective of graph data management, systematically investigating how graph techniques enhance each core stage of the RAG process, as illustrated in Figure 2. Specifically, we highlight how graph data management methods significantly (i) improve knowledge base construction; (ii) optimize retrieval and prompting algorithms through graph-based indexing, querying, and reasoning; (iii) streamline data processing pipelines via graph-structured workflows; and (iv) effectively support graph-oriented tasks.
@@ 4.1.1. A Unified View
In this section, we focus on existing knowledge graphs, which play a crucial role in many RAG systems by serving as structured repositories of factual knowledge. To enable efficient querying and traversal of entities and their relationships, many existing methods (Hu et al., 2024; Baek et al., 2023b; Jiang et al., 2023b) directly retrieve knowledge relevant to the query from well-established knowledge graphs such as FreeBase (Bollacker et al., 2008), T-REx (ElSahar et al., 2018), and WebQSP (Yih et al., 2016). These widely used databases offer comprehensive scopes for knowledge retrieval, encompassing a vast array of entities and relationships across diverse domains. For instance, T-REx (ElSahar et al., 2018) provides extensive alignments between natural language expressions and knowledge base triples, facilitating the integration of structured data with textual information. By organizing nodes and edges along with their textual attributes, these structured repositories allow RAG systems to efficiently access additional topological information, supporting advanced operations such as entity retrieval, relationship discovery, and complex multi-hop queries.
@@ 4.1.2. Existing Knowledge Graphs
Existing KGs Broad coverage, reliable quality General knowledge, less adaptable BioRED (Luo et al., 2022), QALD-9-plus (Perevalov et al., 2022), OpenbookQA (Mihaylov et al., 2018), CREAK (Onoe et al., 2021), TriviaQA (Joshi et al., 2017), HotpotQA (Yang et al., 2018), Mintaka (Sen et al., 2022), MedQA (Jin et al., 2020), TUDataset (Morris et al., 2020), CWQ (Talmor and Berant, 2018), Beyond I.I.D.  (Gu et al., 2021), CommonsenseQA (Talmor et al., 2019), SocialIQA (Sap et al., 2019b), PIQA (Bisk et al., 2020), RiddleSense (Lin et al., 2021), Freebase (Jiang et al., 2019), ATOMIC (Sap et al., 2019a), FactKG (Kim et al., 2023b), MultiHop-RAG (Tang and Yang, 2024), T-REx (ElSahar et al., 2018), DBpedia (Auer et al., 2007a), Yago (Suchanek et al., 2007)
@@
Graphs generated from texts Domain-specific, easily updated LM-dependent, computationally intensive GraphRAG (Edge et al., 2024), GRBK (Delile et al., 2024), ATLANTIC (Munikoti et al., 2023), GNN-Ret (Li et al., 2024f), HippoRAG (Gutiérrez et al., 2024), DALK (Li et al., 2024a), KGP (Wang et al., 2024d), OpenSCR (Han et al., 2023), MindMap (Wen et al., 2024), FABULA (Ranade and Joshi, 2023), GER (Wu et al., 2023a), FoodGPT (Qi et al., 2023), ChatKBQA (Luo et al., 2024a), MultiQG (Lan and Jiang, 2020), HSGE (Sun et al., 2023), ReTraCk (Chen et al., 2021), RNG-KBQA (Ye et al., 2021), ArcaneQA (Gu and Su, 2022), HybridRAG (Sarmah et al., 2024), EWEK-QA (Dehghan et al., 2024), KG-FiD (Yu et al., 2022), REANO (Fang et al., 2024), MedGraphRAG (Wu et al., 2024b), MINERVA (Das et al., 2018)
@@
Freebase (Bollacker et al., 2008) is a large-scale, structured knowledge database designed to organize and store general human knowledge. It combines the efficiency and scalability of structured databases with the rich semantic representation found in sources like Wikipedia (Bridge, 2001), enabling flexible integration of diverse knowledge. By structuring information into tuples, Freebase facilitates efficient querying, entity linking, and knowledge retrieval across multiple domains. Freebase features an HTTP-based API powered by the Metaweb Query Language (MQL) (Flanagan, 2007), which facilitates intuitive object-oriented queries and schema evolution. Its complete normalization philosophy ensures unique global identifiers (GUIDs) for real-world entities, while its lightweight typing system supports multiple data. In the context of RAG, Freebase serves as a powerful auxiliary database, enabling methods to retrieve structured knowledge for enhanced reasoning and answer generation. With its ability to handle vast datasets and diverse relationships, Freebase provides a robust foundation for integrating external knowledge into RAG systems, improving their capability to process complex queries and generate context-aware responses.
@@
T-REx (ElSahar et al., 2018) is a large-scale dataset designed to align natural language from Wikipedia abstracts with knowledge base (KB) triples from Wikidata. Addressing the alignment problem, T-REx connects unstructured text with structured knowledge representations by providing 11 million triples aligned with over 3 million Wikipedia abstracts, covering more than 600 unique predicates. Its customizable alignment pipeline incorporates techniques like predicate linking, co-reference resolution, and distant supervision to create high-quality alignments. Compared to existing datasets, T-REx provides a larger scale, broader predicate coverage, and high accuracy. This dataset is instrumental in advancing tasks like relation extraction, KB population, and question answering, and serves as a foundational resource for RAG systems to retrieve structured knowledge and improve alignment for complex reasoning.
Figure 3. Timeline of recent graph database
4.1.3. Industrial Graph Database for LLM
Recently, widely used graph databases have moved from early research backends to platforms that support LLM use cases with built-in embedding search, standard query languages, and managed GraphRAG pipelines. As shown in Figure 3, we list these important databases on a timeline: Wikidata’s public query service as a common open graph backend [^1] https://query.wikidata.org/; Amazon Neptune as a managed graph service, later adding Neptune Analytics for large-scale graph and embedding analytics [^2] https://aws.amazon.com/cn/neptune/; Neo4j with mature property-graph capabilities and native vector indexes for hybrid retrieval [^3] https://neo4j.com/release/neo4j-4-0/; JanusGraph as common open-source choices for large or real-time workloads [^4] https://janusgraph.org/; ArangoDB as a multi-model system that added vector indexes [^5] https://www.arangodb.com/docs/3.12/release-notes/; vendors beginning to productize GraphRAG (for example, NebulaGraph; Bedrock Knowledge Bases on Neptune) [^6] https://www.nebula-graph.io/posts/announcing_nebulagraph_RAG; and the GQL graph query language becoming an ISO standard that aligns vendor ecosystems [^7] https://www.iso.org/standard/76120.html.
4.1.4. Trend of Graph Database Construction
In the LLM era, graph databases have evolved along four directions that directly address freshness, cost of construction, and LLM integration. First, standardization has begun to reduce vendor lock-in and smooth model portability. For example, the publication of graph query language (e.g., ISO/IEC 39075:2024  [^8] https://opencypher.org/, https://www.iso.org/standard/76120.html, which formalizes a property-graph data and query language and enables more consistent schemas, queries, and tooling across systems. Second, hybrid graph–vector indexing has become more popular so that semantic similarity can be combined with graph traversal. For instance, Neo4j made vector indexes generally available and documents RAG workflows on top of them  [^9] https://neo4j.com/docs/cypher-manual/5/indexes/semantic-indexes/vector-indexes/, while ArangoDB introduced native vector indexes and vector search functions of its declarative language  [^10] https://docs.arangodb.com/stable/. Third, to lower update latency and construction overhead, change-data-capture (CDC) and streaming ingestion became first-class, such as Neo4j’s CDC feature and Kafka connector strategies that propagate graph updates as events, reducing the gap between source changes and graph availability for RAG/agent pipelines  [^11] https://neo4j.com/blog/developer/change-data-capture-cdc-ga/. Finally, the database construction increasingly blends general encyclopedic KGs with domain KGs for higher-quality grounding in graph-based RAG systems, a trend visible in community benchmarks such as GR-Bench (Jin et al., 2024b) and GraphQA (He et al., 2024). In summary, these updates reduce data-to-serving latency and integration cost, making graph databases easier to operate at scale while preserving the multi-hop structure and provenance required by LLM applications.